# Data-Juicer Sandbox: A Feedback-Driven Suite for Multimodal Data-Model Co-development

### v2

@@ Abstract
The emergence of multimodal large models has advanced artificial intelligence, introducing unprecedented levels of performance and functionality. However, optimizing these models remains challenging due to historically isolated paths of model-centric and data-centric developments, leading to suboptimal outcomes and inefficient resource utilization. In response, we present a new sandbox suite tailored for integrated data-model co-development. This sandbox provides a feedback-driven experimental platform, enabling cost-effective iteration and guided refinement of both data and models. Our proposed “Probe-Analyze-Refine” workflow, validated through practical use cases on multimodal tasks such as image-text pre-training with CLIP, image-to-text generation with LLaVA-like models, and text-to-video generation with DiT-based models, yields transferable and notable performance boosts, such as topping the VBench leaderboard. Extensive experiments also uncover fruitful insights into the interplay between data quality, diversity, model behavior, and computational costs. All codes, datasets, and models are open-sourced to foster future research and applications that would otherwise be infeasible due to the lack of a dedicated co-development infrastructure.
Machine Learning, ICML
@@ 1 Introduction
However, the development trajectories of these models and datasets have historically diverged, guided more by intuition than by systematic co-development methodologies. Recent advances in enhancing multimodal large models tend to be either model-centric or data-centric, rarely bridging these two aspects closely. For example, model-centric methods focus on algorithmic enhancements and architectural innovations under fixed data priors, while data-centric strategies usually focus on processing datasets independently of specific model training contexts (Qin et al., 2024). Both approaches usually suffer from a lack of systematic principles and cooperative synergy, relying heavily on heuristic exploration and single-perspective expertise. This fragmented landscape presents a notable barrier to achieving optimal performance, as the interplay between data characteristics and model capabilities remains largely under-exploited.
@@
To fill this gap, we introduce the Data-Juicer Sandbox, a feedback-driven suite for facilitating the co-development of multimodal data and models. Building upon an open-source data processing system tailored for multimodal large models, Data-Juicer (Chen et al., 2024a, b), our sandbox suite integrates a wealth of off-the-shelf components optimized for usability and compatibility with open-source model-centric infrastructures. Collectively, it offers customizable orchestration from different levels including end-to-end workflows, specific development behaviors, and underlying data-model development capabilities. Within the sandbox laboratory, users are empowered to optimize data and models based on their fruitful feedback in cost-controlled environment. This accelerates insight discovery and informed decision-making, paving the way for transferable, resource-efficient data-model optimization.
To exemplify the efficacy of the sandbox, we propose a “Probe-Analyze-Refine” workflow, crafted to explore the interplay between data processing operators (OPs), model performance feedback, and the transferability of these enhancements. We apply this workflow to four cutting-edge models: Mini-Gemini (Li et al., 2024b), an LLaVA-inspired model for image-to-text generation, EasyAnimate (Xu et al., 2024b) and T2V-Turbo (Li et al., 2024a), two Diffusion Transformer based models for text-to-video generation, and a CLIP model (Gadre et al., 2023) for image-text pre-training. Our extensive experiments show notable evidence of the Sandbox’s usability, including topping the VBench (Huang et al., 2024) leaderboard with superior performance than competitors such as Gen-3 (RunwayML, 2024) and VEnhancer (He et al., 2024a). The improvement is underpinned by a series of insights linking more than 40 data processing OPs and 70 benchmark metrics, with fine-grained analysis on the balance between data quality, diversity, compute cost, and model performance.
@@
  • To the best of our knowledge, this is the first open-source  [^1] https://github.com/modelscope/data-juicer/blob/main/docs/Sandbox.md sandbox suite tailored for co-development between multimodal data and models, rendering experimental exploration in this field more insightful and systematic.
@@ 2 Related Works
Open-Source Infrastructure. The ecosystem for multimodal model development has expanded with fruitful open-sourced frameworks (Wolf et al., 2020; Liu et al., 2023). However, contributions to multimodal data infrastructure often consist of raw datasets and preprocessing scripts, lacking standardized practices. Most existing data processing frameworks are tailored for single-modal data (TogetherAI, 2023), highlighting the early stage of multimodal data development. To address these gaps, our work presents a new intermediary layer that connects advanced model infrastructures with the Data-Juicer system, facilitating better co-development between models and data.
Detailed comparison on more works are in Appendix A.
Figure 1: A probe-analyze-refine workflow within Data-Juicer sandbox for systematic data-model co-development. (1) The process begins with cost-effective, contrastive explorations on small data pools to identify the most effective data processing operators (OPs). (2) These OPs are then combined and scaled within data recipes through hierarchical data pyramids to assess their cumulative impact on model performance. (3) Lastly, data utilization is optimized via duplication and diversity analysis, guiding the final formulation of refined data recipes and datasets for enhanced large-scale training. This structured workflow enables informed decision-making and efficient optimization for developing superior models.
@@ 3.1 Motivation and Overview
Why do we need data-model co-development? In the era of large models, the development of both data and models necessitates collaboration involving numerous algorithm researchers and system engineers. Training data for large models is often highly heterogeneous in terms of quality, context, type, and timeliness. The processing and mixing of these datasets, known as data recipes, are complex and varied (Ge et al., 2024). The scale of data amplifies the stakes for refinement attempts on both data and models, imposing substantial computational and time burdens. Traditional data-centric or model-centric strategies fall short by optimizing in isolation, leading to diminished overall efficiency and resource misallocation (Qin et al., 2024). When either component requires adjustment, the dual optimization challenge inflates costs, as one part may have already reached near-optimal status.
@@
Why do we need a sandbox laboratory? Given the high cost associated with iterative development, existing methods often resort to heuristic approaches for improving data or models. For example, scaling up “cleaned” datasets can be problematic because determining what constitutes a “clean” dataset and measuring its quality qualitatively remains a challenge. Further, the impact of iterations on data and models is difficult to attribute due to numerous influencing factors and considerable engineering effort required. More insightful solutions are thus needed for data-model co-development. A unified sandbox environment shows promising, with high potentials to enable experimentation with guided optimization. It permits users to iterate both data and models based on comparative feedback from small-scale data processing and model training trials, thus helping to mine insights that can be transferred to full-scale production environments and increasing the return on investment in computation and time costs.
@@
Layered orchestration of Data-Juicer sandbox laboratory. Due to the current absence of ready-made open-source middleware infrastructure, we first propose the necessary codebase and suite, designed with decoupled components. They serves for data analysis, processing, recipe optimization, model training, and evaluation, all managed through configuration files and divided into three levels: (1) In top-level, the specific workflows are formed as ordered job list for co-development, organized into four phases—probing data/models, refining data recipes, executing data/model operations, and evaluation. One can flexibly adjust task sequences in these phases, reuse built-in workflows or create their own easily. (2) In middle-level, the common behaviors are formed as hook functions; and (3) in the bottom layer, the necessary capabilities are formed as factory classes. The lower two layers work together to facilitate actionable development capabilities and provide measurable feedback. Specifically, we simplify the interfaces offered by Data-Juicer, which provides over 100 OPs and tools for multimodal data analysis, filtering, and synthesis, with industry-level optimization for large-scale processing (Chen et al., 2024b). Built upon it, we integrate many SOTA open-source libraries for model training and evaluation such as Mini-Gemini, EasyAnimate, ModelScope, VBench, MMBench (Liu et al., 2023), TextVQA (Singh et al., 2019) and MME (Fu et al., 2023).
@@
Note that the underlying libraries are still evolving, and the sandbox is continuously maintained to incorporate more community efforts in other libraries, optimized for usability and reduced cognitive load. Dedicated scripts combine the classes and hooks into specific workflows, with code and parameters adjusted to ensure a cost-controlled environment. This suite serves as a groundwork for deriving following experimental insights. More details from the infrastructure perspective are in Appendix B.
@@ 3.2 A Probe-Analyze-Refine Workflow
(3) Built upon found OP combinations, we delve into a dual analysis focusing on data duplication and diversity (Sec. 3.2.3). We assess whether the model training would benefit more compute from the repeated use of high-quality data pools or from the inclusion of lower-quality data to expand the overall data pool. As a result, we obtain data recipes optimized from numerous small-scale comparative experiments, which can then be applied to larger-scale scenarios with reduced amortized cost (Sec 3.2.4).
@@ 3.2.1 Single-Operator Data Pools
Subsequently, models are trained independently on $\{\mathcal{P}_{i}\}$ and $\mathcal{D}_{rand}$ with consistent hyper-parameters, compute, and evaluated across interested performance metrics. This linking of feedback signals between data stats and model metrics enables insight mining and top OP identification.
@@ 3.2.2 Multi-Operators Data Pools
The number of possible $\mathcal{S}$ grows exponentially with $n_{s}$, necessitating non-trivial selection of combinations to explore. Building on insights from single-OP experiments, where the “Top” OPs are represented by their ranks, we propose two practical strategies: (1) combining “Top” OPs based on their progressively diminishing impacts on model performance, and (2) clustering OPs based on their Pearson correlation coefficients and then combining the “Top” OPs within each category. Similar to the single-OP scenario, we then consistently train reference models on $\{\mathcal{P}_{\mathcal{S}}\}$ and $\mathcal{D}_{rand}$ to gain comparative outcomes. More implementation details and empirical results on these two strategies can be found in Appendix D.1 and Sec. 4.3 respectively.
@@ 3.2.4 Discussion on Scaling and Cost
Achieving $(1+mr)\leq M$ is feasible. Typically, $M>2$ accounting for multiple iterations over different model and dataset versions. The $m$ positively corresponds to the interested OPs, often numbering in the dozens. The $r$ can be much smaller than 0.01, since there is no need for small-pool experiments to “converge” as long as they reveal informative changes—positive or negative—in model performance after controlled data interventions with $\{\mathcal{P}_{i},\mathcal{P}_{S}\}$ versus baselines using $\mathcal{D}_{rand}$. Early stopping of unpromising experiment trials can also help expedite this process.
@@ 4.1 Use-Cases and Experiment Overview
To demonstrate the proposed suite’s usability and effectiveness, aiding understanding and boosting confidence in its applicability, in this section, we conduct extensive experiments in three practical use cases. They vary in models and data OPs (Sec. 4.2 and Sec 4.3), data samples and compute cost (Sec. 4.4), leading to insights and recipes useful for larger-scale scenarios (Sec. 4.5).
Specifically, for image-to-text (I2T) generation, we will show that the optimal recipe derived from small pools achieves superior model performance and higher data efficiency on the MGM model when applied to larger datasets. For text-to-video (T2V) generation, we will show that the insights derived from small pools and the EasyAnimate model can lead to VBench Top-1 performance with another architecture-different model. For image-text pre-training (ITP), we will show that the optimal recipe identified with a CLIP model having fewer FLOPS maintains consistent performance advantages when model FLOPS and compute resources increase, similar to scaling law behaviors. These examples instantiate different behavior hooks and capability classes introduced in Sec. 3.1, following the workflow proposed in Sec. 3.2. A more structural summarization of these use cases is provided in Table 5 in Appendix E.1.
In the next subsections, we delve into the primary insights, with complete experimental results provided in Appendix E. The feedback utilized in these experiments comes from over 70 widely-used model benchmark measurements and over 40 Data-Juicer’s data filtering OPs for text, image, video, and cross-modalities. For brevity and comprehensive investigation, in the main page, we report the performance changes relative to baselines on $\mathcal{D}_{rand}$, averaging from all evaluated benchmark scores. For fair comparison, note that all experiments are maintained to ensure that the reference models trained on $\mathcal{P}_{i}$ and $\mathcal{P}_{\mathcal{S}}$ have the same compute costs to baselines trained on $\mathcal{D}_{rand}$. More detailed task-specific settings can be found in Appendix D, including the benchmark measurements (in Appendix D.2), the functionalities of studied OPs (in Appendix D.3), and the specific sources and sizes of ($\mathcal{D},\mathcal{D}_{rand},\mathcal{P}_{i},\mathcal{P}_{\mathcal{S}}$), in Appendix D.4 to D.6.
Table 1: The average performance changes relative to baseline models on $\mathcal{D}_{rand}$ for top-3 performing OPs, with models trained on different single-OP data pools split by sorting OP-generated statistics: $\mathcal{P}_{i,low},\mathcal{P}_{i,mid},\mathcal{P}_{i,low}$.
@@
width = colspec = Q[55]Q[510]Q[113]Q[113]Q[113], cell11 = r=2, cell12 = r=2, cell13 = c=30.4, cell31 = r=3, cell61 = r=3, cell91 = r=3, hline1,12 = -0.08em, hline2 = 3-5, hline3,6,9 = -, Task & OP Statistics Avg. Perf. Changes (%)

### v3

@@ Abstract
The emergence of multimodal large models has advanced artificial intelligence, introducing unprecedented levels of performance and functionality. However, optimizing these models remains challenging due to historically isolated paths of model-centric and data-centric developments, leading to suboptimal outcomes and inefficient resource utilization. In response, we present a new sandbox suite tailored for integrated data-model co-development. This sandbox provides a feedback-driven experimental platform, enabling cost-effective iteration and guided refinement of both data and models. Our proposed “Probe-Analyze-Refine” workflow, validated through practical use cases on multimodal tasks such as image-text pre-training with CLIP, image-to-text generation with LLaVA-like models, and text-to-video generation with DiT-based models, yields transferable and notable performance boosts, such as topping the VBench leaderboard. A comprehensive set of over 100 experiments demonstrated the suite’s usability and extensibility, while also uncovering insights into the interplay between data quality, diversity, model behavior, and computational costs. All codes, datasets, and models are open-sourced to foster future research and applications that would otherwise be infeasible due to the lack of a dedicated co-development infrastructure.
Data Processing, Foundation Models, Sandbox
@@ 1 Introduction
However, the development trajectories of these models and datasets have historically diverged, guided more by intuition than by systematic co-development methodologies. Recent advances in enhancing multimodal large models tend to be either model-centric or data-centric, rarely bridging these two aspects closely. For example, model-centric methods focus on algorithmic enhancements and architectural innovations under fixed data priors, while data-centric strategies usually focus on processing datasets independently of specific model training contexts (Qin et al., 2025). Both approaches usually suffer from a lack of systematic principles and cooperative synergy, relying heavily on heuristic exploration and single-perspective expertise. This fragmented landscape presents a notable barrier to achieving optimal performance, as the interplay between data characteristics and model capabilities remains largely under-exploited.
@@
To fill this gap, we introduce the Data-Juicer Sandbox, a feedback-driven suite for facilitating the co-development of multimodal data and models. Building upon an open-source data processing system tailored for multimodal large models, Data-Juicer (Chen et al., 2024a, b), our sandbox suite integrates a wealth of off-the-shelf components optimized for usability and compatibility with open-source model-centric infrastructures. Collectively, it offers customizable orchestration from different levels including end-to-end workflows, specific development behaviors, and underlying data-model development capabilities. Within the sandbox laboratory, users are empowered to optimize data and models based on their fruitful feedback in a cost-controlled environment. This accelerates insight discovery and informed decision-making, paving the way for transferable, resource-efficient data-model optimization.
To exemplify the efficacy of the sandbox, we propose a “Probe-Analyze-Refine” workflow, crafted to explore the interplay between data processing operators (OPs), model performance feedback, and the transferability of these enhancements. We apply this workflow to five cutting-edge models: Mini-Gemini (Li et al., 2024b) and InternVL-2.0 (Chen et al., 2024f), two LLaVA-inspired models for image-to-text generation, EasyAnimate (Xu et al., 2024b) and T2V-Turbo (Li et al., 2024a), two Diffusion Transformer based models for text-to-video generation, and a CLIP model (Gadre et al., 2023) for image-text pre-training. Through over 100 experimental runs and comparative analysis across different tasks, scales and models, we show notable evidence of the Sandbox’s usability, including topping the VBench (Huang et al., 2024) leaderboard with superior performance over competitors such as Gen-3 (RunwayML, 2024) and VEnhancer (He et al., 2024a). The improvement is underpinned by a series of insights linking more than 40 data processing OPs and 70 benchmark metrics, with fine-grained analysis on the balance between data quality, diversity, compute cost, and model performance, such as scaling behaviors derived from 1B to 26B mode scales.
@@
  • To the best of our knowledge, this is the first open-source 111modelscope.github.io/data-juicer/en/main/docs/Sandbox.html sandbox suite tailored for co-development between multimodal data and models, rendering experimental exploration in this field more insightful and systematic.
@@ 2 Related Works
Open-Source Infrastructure. The ecosystem for multimodal model development has expanded with fruitful open-sourced frameworks (Wolf et al., 2020; Liu et al., 2024b). However, contributions to multimodal data infrastructure often consist of raw datasets and preprocessing scripts, lacking standardized practices. Most existing data processing frameworks are tailored for single-modal data (Weber et al., 2024), highlighting the early stage of multimodal data development. To address these gaps, our work presents a new intermediary layer that connects advanced model infrastructures with the Data-Juicer system, facilitating better co-development between models and data.
We present more detailed comparisons of more related works in Appendix A.
Figure 1: A probe-analyze-refine workflow within Data-Juicer sandbox for systematic data-model co-development.
@@ 3.1 Motivation and Overview
Why do we need data-model co-development? In the era of large models, the development of both data and models necessitates collaboration involving numerous algorithm researchers and system engineers. Training data for large models is often highly heterogeneous in terms of quality, context, type, and timeliness. The processing and mixing of these datasets, known as data recipes, are complex and varied (Ge et al., 2024). The scale of data amplifies the stakes for refinement attempts on both data and models, imposing substantial computational and time burdens. Traditional data-centric or model-centric strategies fall short by optimizing in isolation, leading to diminished overall efficiency and resource misallocation (Qin et al., 2025). When either component requires adjustment, the dual optimization challenge inflates costs, as one part may have already reached near-optimal status.
@@
Why do we need a sandbox laboratory? Given the high cost associated with iterative development, existing methods often resort to heuristic approaches for improving data or models. For example, scaling up “cleaned” datasets can be problematic because determining what constitutes a “clean” dataset and measuring its quality qualitatively remains a challenge. Further, the impact of iterations on data and models is difficult to attribute due to numerous influencing factors and considerable engineering effort required. A unified sandbox environment enables experimentation with guided optimization. It permits users to iterate both data and models based on comparative feedback from small-scale data processing and model training trials, thus helping to mine insights that can be transferred to full-scale production environments and increasing the return on investment in computation and time costs.
@@
Layered orchestration of Data-Juicer sandbox laboratory. Due to the current absence of ready-made open-source middleware infrastructure, we first propose the necessary codebase and suite, designed with decoupled components. They serve for data analysis, processing, recipe optimization, model training, and evaluation, all managed through configuration files and divided into three levels: (1) In the top level, the specific workflows are formed as ordered job list for co-development, organized into four phases—probing data/models, refining data recipes, executing data/model operations, and evaluation. One can flexibly adjust task sequences in these phases, reuse built-in workflows, or create their own easily. (2) In the middle level, the common behaviors are formed as hook functions; and (3) in the bottom layer, the necessary capabilities are formed as factory classes. The lower two layers work together to facilitate actionable development capabilities and provide measurable feedback. Specifically, we simplify the interfaces offered by Data-Juicer, which provides over 100 OPs and tools for multimodal data analysis, filtering, and synthesis, with industry-level optimization for large-scale processing (Chen et al., 2024b). Built upon it, we integrate many SOTA open-source libraries for model training and evaluation such as Mini-Gemini, EasyAnimate, ModelScope, VBench, MMBench (Liu et al., 2024b), TextVQA (Singh et al., 2019) and MME (Fu et al., 2023).
@@
Note that the underlying libraries are still evolving, and the sandbox is continuously maintained to incorporate more community efforts in other libraries, optimized for usability and reduced cognitive load. Dedicated scripts combine the classes and hooks into specific workflows, with code and parameters adjusted to ensure a cost-controlled environment. This suite serves as a groundwork for the deriving following experimental insights. More details from the infrastructure perspective are in Appendix B.
@@ 3.2 A Probe-Analyze-Refine Workflow
(3) Built upon found OP combinations, we delve into a dual analysis focusing on data duplication and diversity (Sec. 3.2.3). We assess whether the model training would benefit more compute from the repeated use of high-quality data pools or from the inclusion of lower-quality data to expand the overall data pool. As a result, we obtain data recipes optimized from numerous small-scale comparative experiments, which can then be applied to larger-scale scenarios with reduced amortized cost (Sec. 3.2.4).
@@ 3.2.1 Single-Operator Data Pools
Subsequently, models are trained independently on $\{\mathcal{P}_{i}\}$ and $\mathcal{D}_{rand}$ with consistent hyper-parameters, data sample size and compute resources, and finally evaluated across interested performance metrics. This linking of feedback signals between data stats and model metrics enables insight mining and top OP identification.
@@ 3.2.2 Multi-Operators Data Pools
The number of possible $\mathcal{S}$ grows exponentially with $n_{s}$, necessitating non-trivial selections of combinations to explore. Building on insights from single-OP experiments, where the “Top” OPs are represented by their ranks, we propose two practical strategies: (1) combining “Top” OPs based on their progressively diminishing impacts on model performance, and (2) clustering OPs based on their Pearson correlation coefficients and then combining the “Top” OPs within each category. Similar to the single-OP scenario, we then consistently train reference models on $\{\mathcal{P}_{\mathcal{S}}\}$ and $\mathcal{D}_{rand}$ to gain comparative outcomes. More implementation details and empirical results on these two strategies can be found in Appendix D.1 and Sec. 4.3 respectively.
@@ 3.2.4 Discussion on Scaling and Cost
Achieving $(1+mr)\leq M$ is feasible. Typically, $M>2$ accounts for multiple iterations over different model and dataset versions. The $m$ positively corresponds to the interested OPs, often numbering in the dozens. The $r$ can be much smaller than 0.01, since there is no need for small-pool experiments to “converge” as long as they reveal informative changes—positive or negative—in model performance after controlled data interventions with $\{\mathcal{P}_{i},\mathcal{P}_{S}\}$ versus baselines using $\mathcal{D}_{rand}$. Early stopping of unpromising experiment trials can also help expedite this process.
@@ 4.1 Use-Cases and Experiment Overview
To demonstrate the proposed suite’s usability and effectiveness, aiding understanding and boosting confidence in its applicability, in this section, we conduct extensive experiments in various practical use cases. They vary in models and data OPs (Sec. 4.2 and Sec. 4.3), data samples and compute cost (Sec. 4.4), leading to insights and recipes useful for larger-scale scenarios (Sec. 4.5).
  • For image-to-text (I2T) generation, we will show that the optimal recipe derived from small pools achieves superior model performance and higher data efficiency on the MGM model when applied to larger datasets.
  • For general text-to-video (T2V) generation, we will show that the insights derived from small pools and the EasyAnimate model can lead to VBench Top-1 performance with another architecture-different model.
  • For image-text pre-training (ITP), we will show that the optimal recipe identified with a CLIP model having fewer FLOPS maintains consistent performance advantages when model FLOPS and compute resources increase, similar to scaling law behaviors.
  • For image-captioning (IC) oriented fine-tuning, we will show that the optimal recipe identified with InternVL-2.0 model (Chen et al., 2024f), leads to consistent and steady performance advantages as the model scale increases from 1B to 26B.
  • To demonstrate the flexibility, extensibility, and potential of the proposed sandbox, we further conduct experiments on “iterative workflow”, “OPs beyond filters”, and “model development of automated prompt adjustments”. More details will be introduced later in Sec. 4.6.
These examples instantiate different behavior hooks and capability classes introduced in Sec. 3.1, following the workflow proposed in Sec. 3.2. A more structural summarization of these use cases is provided in Table 5 in Appendix E.1.
In the next subsections, we delve into the primary insights, with complete experimental results provided in Appendix E. The feedback utilized in these experiments comes from over 70 widely-used model benchmark measurements and over 40 Data-Juicer’s data filtering OPs for text, image, video, and cross-modalities. For brevity and comprehensive investigation, on the main page, we report the performance changes relative to baselines on $\mathcal{D}_{rand}$, averaging from all evaluated benchmark scores. For a fair comparison, note that all experiments are maintained to ensure that the reference models trained on $\mathcal{P}_{i}$ and $\mathcal{P}_{\mathcal{S}}$ have the same compute costs to baselines trained on $\mathcal{D}_{rand}$. Detailed task-specific settings can be found in Appendix D, including the benchmark measurements (in Appendix D.2), the functionalities of studied OPs (in Appendix D.3), and the specific sources and sizes of ($\mathcal{D},\mathcal{D}_{rand},\mathcal{P}_{i},\mathcal{P}_{\mathcal{S}}$), in Appendix D.4 to D.6.
Table 1: The average performance changes relative to baseline models on $\mathcal{D}_{rand}$ for top-3 performing OPs, with models trained on different single-OP data pools split by sorting OP-generated statistics: $\mathcal{P}_{i,low},\mathcal{P}_{i,mid},\mathcal{P}_{i,low}$. The training data pool for baseline is randomly sampled, and all compared pools are with equal data volume. Full ranking table can be found in Appendix E.2.
@@
width = colspec = Q[55]Q[510]Q[113]Q[113]Q[113], cell11 = r=2, cell12 = r=2, cell13 = c=30.4, cell31 = r=3, cell61 = r=3, cell91 = r=3, cell121 = r=3, hline1,15 = -0.08em, hline2 = 3-5, hline3,6,9,12 = -, Task & OP Statistics Avg. Perf. Changes (%)
@@
IC Text Length 0.76 -3.13 -11.36
Image Watermark Score -0.64 -0.13 0.48
Character Repetition Ratio 0.45 -0.46 -0.63